OpenStudy (anonymous):

find the derivative: y=(4xe^x)/(x^2+1)

OpenStudy (anonymous):

u = 4xe^x v = x^2 + 1 u' = 4(e^x + xe^x) v' = 2x (u'v -v'u) / v^2 (4(e^x+xe^x)*(x^2+1) - 2x*4xe^x) / (x^2 +1)^2

OpenStudy (anonymous):

that seems very long to be the answer...

OpenStudy (anonymous):

you have to solve the algebra

OpenStudy (anonymous):

what algebra?

OpenStudy (anonymous):

simplify it you see there are multiplications there

myininaya (myininaya):

\[\ln(y)=\ln(\frac{4xe^{x}}{x^2+1})\] \[\ln(y)=\ln(4xe^x)-\ln(x^2+1)\] \[\ln(y)=\ln(4)+\ln(x)+\ln(e^x)-\ln(x^2+1)\] \[\ln(y)=\ln(4)+\ln(x)+x \ln(e)-\ln(x^2+1)\] \[\ln(y)=\ln(4)+\ln(x)+x-\ln(x^2+1)\] \[\frac{y'}{y}=0+\frac{1}{x}+1-\frac{2x}{x^2+1}\] \[y'=y(0+\frac{1}{x}+1-\frac{2x}{x^2+1})\] \[y'=\frac{4xe^{x}}{x^2+1} \cdot(1+\frac{1}{x}-\frac{2x}{x^2+1})\]

OpenStudy (anonymous):

is that the final answer?